AAVSOnet status, 2016-04-21

Thu, 04/21/2016 - 12:15

We've stopped image processing for a while, as we are debugging some software.  Data continues to be collected, so those of you with research projects will not be affected other than a delay before you can reach your images.

BSM South is now working with its normal ST-8XME camera.

BSM_NM is nearly back on-line.  Bill Stein replaced the CGEM mount with our spare Orion EQ-6 mount, and Aaron Sliski is working with that mount to have it interface properly with the acquisition software.  We polar aligned the telescope, and Bill and Aaron adjusted the focuser so that it moves reliably.  We expect to be taking survey data shortly.

The mount for BSM Argentina has been repaired and will be installed shortly.  We are also sending down a webcam so that we can watch the telescope remotely.

BSM Berry is still down for mount repairs.

BSM HQ is working reasonably well, though it has an intermittent communication issue with the mount.

BSM Hamren's computer failed.  We wanted to replace it with a lower-power solution anyway, as this system is "off the grid".  Mike Linnolt has sent the hard drive back to HQ so that we can offload the last few nights of data, and we hope to have the alternative computer solution working within a week or two.

Coker30 is running nominally.

SRO is running nominally.

OC61 is running nominally.  Nigel Frost has the focal reducer in-hand for the spectrographic channel, and is now working on installing it.

TMO61's dome slit has been repaired.  Jon Holtzman is working with a student to test the LISA spectrograph for the next few days, but we hope to return to normal imaging once the moon fades a bit.